OpenStudy (anonymous):

Use substitution for this problem. 3x+y=5 6x=10-2y

OpenStudy (anonymous):

I already have like... 3x+y=5 6x+2y=10

Vocaloid (vocaloid):

hm, it wants you to use substitution start from the beginning

Vocaloid (vocaloid):

3x+y=5 6x=10-2y now divide the second equation by 2

OpenStudy (anonymous):

The whole equation

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

yes, divide the second equation by 2

OpenStudy (anonymous):

i got 3x+y=5 3x+y=5

Vocaloid (vocaloid):

good, so, notice anything special about the two equations?

OpenStudy (anonymous):

they are the same!

Vocaloid (vocaloid):

yeah, so what is our solution?

OpenStudy (anonymous):

(0,0)

Vocaloid (vocaloid):

hmmmm not quite

Vocaloid (vocaloid):

when we have two equations that are the same, that means that there are an infinite number of solutions

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

thank you